OpenStudy (anonymous):

The variables x and y vary directly. Write an equation that relates x and y, when x = 4 and y = 2.

OpenStudy (anonymous):

Varies directly means y = kx where k is some constant. so if y = 2 and x = 4 then 2 = k(4) solve for k 1/2 = k Equation then is y = 1/2x

OpenStudy (anonymous):

thanks